# AED: Adaptable Error Detection for Few-shot Imitation Policy

### v1

@@ Appendix A FSI Policy Details
We follow the main paper’s Section 3 to implement three state-of-the-art demonstration-conditioned (DC) policies, including NaiveDC(James et al., 2018), DCT(Dance et al., 2021), and SCAN(Yeh et al., 2022), and monitor their behaviors.
@@
In the visual head, we added a fully-connected and a dropout layer after a single ResNet18 to avoid overfitting the data from base environments. Furthermore, the input dimension of the actor module depends on the sum of dimensions of the outputs from the visual head and the task-embedding network. For the DCT and SCAN policies, the task-embedding’s dimension matches the visual head’s output. However, the task-embedding’s dimension in NaiveDC policy is twice the size.
Regarding the task-embedding network, NaiveDC concatenates the first and last frames of each demonstration and computes the average as the task-embedding, without involving any attention mechanism. In addition, DCT incorporates cross-demo attention to fuse demonstration features and then applies rollout-demo attention to compute the task-embedding. SCAN uses an attention mechanism to attend to each demonstration from the rollout, filtering out uninformative frames, and fuses them to obtain the final task-embedding. Furthermore, we use a standard LSTM, which differs from Yeh et al. (2022)’s setting.
@@
To optimize the policies, we utilize a RMSProp optimizer (Hinton, ) with a learning rate of 1e-4 and a L2 regularizer with a weight of 1e-2. Each training epoch involves iterating through all base environments. Within each iteration, we sample five demonstrations and ten rollouts from the sampled base environment. The total number of epochs varies depending on the specific tasks and is specified in Table 4. All policy experiments are conducted on a Ubuntu 20.04 machine equipped with an Intel i9-9900K CPU, 64GB RAM, and a NVIDIA RTX 3090 24GB GPU.
@@ Appendix B PrObe Details
In this work, we propose the Pattern Observer (PrObe) to address the novel AED task by detecting behavior errors from the policy’s feature representations, which offers two advantages: (1) PrObe has a better understanding of the policy because the trained policy remains the same during both AED training and testing. This consistency enhances its effectiveness in identifying errors. (2) The additional task knowledge from the policy encoder aids in enhancing the error detection performance. We have described PrObe’s components and their usages in the main paper. Here, we introduce more details about the design of model architecture and training objectives.
Architecture design of PrObe
The pattern extractor aims to extract the observable patterns from embeddings in policy features. To achieve this, we can leverage discrete encoding, as used in Hafner et al. (2022). However, the required dimension of discrete embedding might vary and be sensitive to the evaluation tasks. To address this, we leverage an alternative approach that operates in a continuous but sparse space. Besides, for the pattern flow generator, we use a standard LSTM model instead of a modern transformer. This is motivated by the characteristics of the AED task, where policy rollouts are collected sequentially, and adjacent frames contain crucial information when behavior errors occur. Thus, the LSTM model is better suited for addressing the AED task. The contributions of PrObe’s components are verified and shown in Figure 12. From the results, our designed components significantly improve the performance.
@@
With the objective $L_{pat}$, the pattern embeddings are expected to be sparse and easily observable of changes, benefiting the model in distinguishing behavior errors. Next, as previously pointed out, when applying contrastive learning to temporal-sequence data, the time relationship should also be considered, in addition to determining whether the frames are normal or errors for the task. Thus, $L_{tem}$ is a novel temporal-aware triplet loss (Schroff et al., 2015), and the temporal-aware margin $m_{t}$ in $L_{tem}$ will have different effects depending on the relationship between the anchor and the positive/negative sample, as depicted in Figure 7. The $L_{tem}$ can be calculated by:
@@
where $N$ is the number of sampled pairs, and the temporal-aware margin $m_{t}=m*(1.0-\alpha*(d_{ap}-d_{an}))$ will be enlarged or reduced considering the temporal distance of anchor, positive and negative samples. Among them, $m$ is the margin in the original triplet loss, and $d_{ap}$, $d_{an}$ are the clipped temporal distances between the anchor and positive sample and the anchor and negative sample, respectively. With $L_{tem}$, PrObe can efficiently perform contrastive learning without getting confused by blindly pulling temporally distant anchors and positive samples closer. Lastly, the total loss $L_{total}$ is the combination of three objectives:
@@
method name policy training temporal DC-based output type
dependent rollouts information
SVDDED successful rollouts only distance to center
TaskEmbED both probability
LSTMED both probability
DCTED both probability
PrObe (ours) both probability
@@ Appendix C Error Detector Details
  • SVDDED: A modified one-class SVM method (Ruff et al., 2018) trained only with successful rollouts. It relies on measuring the distance between rollout embeddings to the center embedding to detect behavior errors, without considering the temporal information. We average demonstration features as the center embedding and minimize the embedding distance between rollout features and the center during training.
  • TaskEmbED: A single-frame baseline trained with successful and failed rollouts, which is a modification from the NaiveDC policy (James et al., 2018). It concatenates and averages the first and last demonstration frames as the task-embedding. Then, it predicts the behavior errors conditioned on the current frame and task-embedding.
  • LSTMED: A standard LSTM model (Hochreiter & Schmidhuber, 1997) predicts errors based solely on the current rollout. It is trained with successful and failed rollouts with no demonstration knowledge provided. It is expected to excel in identifying errors that occur at similar timings to those observed during training. However, it may struggle to recognize errors that deviate from the demonstrations.
  • DCTED: A baseline with temporal information derived from the DCT policy (Dance et al., 2021). It is trained with successful and failed rollouts and incorporates with demonstration information. One notable distinction between DCTED and our PrObe lies in their policy dependencies. DCTED detects errors using its own encoder, which solely leverages visual information obtained within the novel environment. Conversely, PrObe learns within the policy feature space, incorporating additional task-related knowledge.
@@ Appendix D Evaluation Tasks
To obtain accurate frame-level error labels in the base environments and to create a simulation environment that closely resembles real-world scenarios, we determined that existing robot manipulation benchmarks/tasks (Yu et al., 2019b; James et al., 2020; Fu et al., 2020; Mandlekar et al., 2021; Yeh et al., 2022; Li et al., 2022; Gu et al., 2023) did not meet our requirements. Consequently, seven challenging FSI tasks containing 329 base and 158 novel environments are developed. Their general attributes are introduced in Section D.1 and Table 4; the detailed task descriptions and visualizations are provided in Section D.2 and Figure 8-9, respectively.
@@ D.1 General Task Attributes
  • Generalization: To present the characteristics of changeable visual signals in real scenes, we build dozens of novel environments for each task, comprising various targets and task-related or environmental distractors. Also, Dozens of base environments are built as training resources, enhancing AED methods’ generalization ability. Letting them train in base environments with different domains can also be regarded as a domain randomization technique (Bonardi et al., 2020), preprocessing for future sim2real experiments.
  • Challenging: Three challenges, including multi-stage tasks, misaligned demonstrations, and different appearances between expert and agent, are included in our FSI tasks. Moreover, unlike existing benchmarks (e.g., RLBench (James et al., 2020)) attaching the objects to the robot while grasping, we turn on the gravity and friction during the whole process of simulation, so grasped objects may drop due to unsmooth movement.
  • Realistic: Our tasks support multi-source lighting, soft shadow, and complex object texture to make them closer to reality (see Figure 9). Existing benchmarks usually have no shadow (or only a hard shadow) and simple texture. The gravity and friction stated above also make our tasks more realistic.
@@ D.2 Task Descriptions
Description
Fully close the drawer that was closed by the expert in demonstrations. Besides, the objects on the table will be randomly set in the designated area.
Success condition
The target drawer is fully closed, and another drawer (distractor) must not be moved.
Behavior errors
(1) Not fully close the target drawer; (2) Close the distractor, not the target drawer; (3) Close two drawers at the same time.
@@
Description
Fully press the button which is randomly placed in a designated area.
Success condition
The button is fully pressed.
Behavior errors
(1) Not touching the button at all; (2) The button is not fully pressed.
@@
Description
Pick and place the mug/cup into the bowl the expert placed in demonstrations.
Success condition
The mug/cup is fully placed in the target bowl.
Behavior errors
(1) Failed to pick up the mug/cup; (2) Failed to place it into the target bowl; (3) Misplaced the mug/cup into another bowl (distractor).
@@
Description
Pick up and place a glass of water on the coaster the expert placed in demonstrations.
Success condition
The glass of water is placed on the target coaster and no water is spilled.
Behavior errors
(1) Failed to pick up the glass of water; (2) drips spilling; (3) Not placed on the target coaster.
@@
Description
Pick up and place the object in front of the target bookshelf, and push it under the shelf.
Success condition
The object is fully under the target shelf.
Behavior errors
(1) Failed to pick up the object; (2) Losing object during moving; (3) Not fully placed under the target shelf.
@@
Description
Pick up the magic cube/dice and place it into the storage box. Then, push the box until it is fully under the bookshelf.
Success condition
The magic cube/dice is in the storage box, and the box is fully under the bookshelf.
Behavior errors
(1) Failed to pick up the magic cube/dice. (2) Failed to place the magic cube/dice into the storage box. (3) The box is not fully under the bookshelf.
@@
Description
Wait until the product box reaches the operating table, place the product in the box, and pick and place the box on the next conveyor belt.
Success condition
The product is inside the box, and the box reaches the destination table.
Behavior errors
(1) Failed to place the product into the box. (2) Failed to pick up the box. (3) Failed to place the box on the next conveyor belt.
@@ Appendix E Additional Experimental Results
E.1 Accuracy Function in Our AED task
Figure 10: Rules of the accuracy function $A(\cdot,\cdot)$ to determine the error prediction results.
As stated in the main paper, the frame-level label is lacking during inference because novel environments only indicate whether the current rollout is a success or failure. Hence, we follow similar rules in Wong et al. (2021) to define the accuracy function $A(\cdot,\cdot)$, i.e., how to determine the error prediction outcome, as elaborated in Figure 10. The $A(\cdot,\cdot)$ determines the prediction result at the sequence-level, which may not sufficiently reflect whether the timing to raise the error is accurate. Thus, we further conduct the timing accuracy experiment (Figure 5 in the main paper).
@@ E.2 Main Experiment’s Details
Table 5: FSI policy performance. Success rate (SR) [$\uparrow$] and its standard deviation (STD) are reported. Notably, STD is calculated across novel environments, rather than multiple experiment rounds. That’s also the reason that some STD values are large, as the difficulty of different novel environments in the same task varies.
Close Drawer Press Button Pick &\&& Place Organize Table
NaiveDC $\mathbf{91.11\pm 20.85\%}$ $51.94\pm 18.79\%$ $55.00\pm 24.98\%$ $12.20\pm 13.93\%$
DCT $80.56\pm 26.71\%$ $\mathbf{80.56\pm 11.65\%}$ $64.05\pm 19.77\%$ $\mathbf{79.00\pm 09.27\%}$
SCAN $88.33\pm 13.94\%$ $75.56\pm 12.68\%$ $\mathbf{71.19\pm 15.38\%}$ $66.60\pm 23.18\%$
Move Glass Cup Back to Box Factory Packing
NaiveDC $39.00\pm 28.63\%$ $08.89\pm 09.21\%$ $45.42\pm 30.92\%$
DCT $51.80\pm 23.99\%$ $29.17\pm 09.46\%$ $\mathbf{88.75\pm 07.11\%}$
SCAN $\mathbf{72.60\pm 11.84\%}$ $\mathbf{58.89\pm 10.87}\%$ $63.75\pm 33.11\%$
@@
The policy performance is originally reported in Figure 4 of the main paper. Here, we provide a comprehensive overview in Table 5. The policy conducts 20 executions (rollouts) for each novel environment to obtain the success rate (SR). Then, the average success rate and its standard deviation (STD) are calculated across those SRs. Some large STD values are due to the variety in difficulty among novel environments. Factors such as variations in object size or challenging destination locations can lead to different performance outcomes for the policies. Additionally, these collected rollouts are also used for later AED inference.
In general, the NaiveDC policy achieves the best results in the simplest tasks because the task-embedding from the concatenation of the first and last two frames provides sufficient information. However, its performance rapidly declines as task difficulty increases. On the other hand, the DCT policy excels in tasks where the misalignment between demonstrations is less severe, as it fuses the task-embeddings in the temporal dimension. Lastly, the SCAN policy outperforms in more difficult tasks because its attention mechanism filters out uninformative demonstration frames.
Main experiment’s detailed values and ROC curves
The main experiment result is originally reported in Figure 4 of the main paper. Here, Table 8 and Figure 15 present its detailed values and ROC curves, which better demonstrate our PrObe’s performance gap against other methods. To highlight again, our proposed PrObe obtains the best performance on both metrics and can handle the various behaviors from different policies.
@@ E.3 Ablation Study
Rollout augmentation
As stated before, the collected agent rollouts exhibit limited diversity. This experiment aims to validate whether rollout augmentation (RA) enhances the rollout diversity and, consequently, improves the robustness of error detectors. In Figure 11, we present the results of all error detection methods (trained with and without RA) monitoring policies solving the Pick &\&& Place task, yielding the following findings: (1) RA has a minor or even negative impact on SVDDED, as expected. Since SVDDED is exclusively trained on (single) frames from successful rollouts, the removal of frames by RA decreases overall frame diversity. (2) RA proves beneficial for methods incorporating temporal information (LSTMED, DCTED, and PrObe), particularly when the FSI policy performs better (SCAN case). This is because error detectors cannot casually trigger an error, or they risk making a false-positive prediction. The results indicate that methods trained with RA generate fewer false-positive predictions, showcasing improved robustness to various policy behaviors.
Contribution of PrObe components
Figure 12 illustrates the performance contribution of each PrObe’s component, resulting from monitoring the SCAN policy solving the Pick & Plcae task. The naive model (first column) removes the pattern extractor and uses a FC layer followed by an IN layer to transform history embeddings. Obviously, the designed components and objectives boost performance, especially when adding the pattern extractor into the model. We attribute this to the fact that extracted embedding (representation) patterns by our components are more informative and discernible.
@@
Table 6: Demonstration quality experiment. AUROC [$\uparrow$] is reported. This experiment verifies the influence of demonstration quality on both the FSI policy and AED methods. PrObe is the only method that can achieve higher performance when receiving sub-optimal demonstrations, as the FSI policies raise more failures.
@@
All optimal $75.56\pm 12.68\%$ 0.4653 0.4657 0.4487 0.6757 0.7505
3 optimal &\&& 2 sub-optimal $71.67\pm 11.18\%$ 0.4985 0.5165 0.4979 0.7467 0.7804
All sub-optimal $69.72\pm 12.30\%$ 0.4886 0.5262 0.5096 0.7140 0.7752
Performance stability
Our work involves the training and testing of all policies and AED methods to produce results for an FSI task. Additionally, each task comprises numerous base and novel environments. These factors make it time-consuming and impractical to include error bars in the main experiment (Figure 4 of the main paper). Based on our experience, generating the results of the main experiment once would require over 150 hours.
Therefore, we selected the most representative Pick &\&& Place task among our FSI tasks and conducted multiple training and inference iterations for all AED methods. We present the AUROC, AUPRC, and their standard deviations averaged over multiple experimental rounds (with five random seeds) in Figure 13. From the results, our proposed PrObe not only achieves the best performance in detecting behavior errors across all policies but also exhibits the most stable performance among all AED methods, with the smallest standard deviation (STD) values.
@@
In this experiment, we aim to verify the influence of demonstration quality on both the FSI policy and AED methods. We test it on the Press Button task. We collect sub-optimal demonstrations that failed to press the button at the first trial but pressed the button successfully on the second try. We provide these sub-optimal demonstrations and let the policies and AED methods perform the task.
Specifically, the policy performance should drops when the demonstration quality decreases. As a result, the AED methods should get a higher AUPRC score if they are not affected by the quality changes of demonstrations since it’s easier to find an error. However, from Table 6, we can observe that not all AED methods obtain higher scores, meaning that they are also affected by the quality decreasing. Notably, our proposed PrObe is the only method that achieves better results for all cases - compare the first and second rows and the first and third rows for each policy. Additionally, an interesting observation can be found in the case of NaiveDC policy (all optimal vs. all sub-optimal): The LSTMED and the TaskEmbED obtained similar AUPRC scores, as they don’t access the demonstration information; SVDDED baseline suffered in this situation since it averages every demonstration frame (including those from sub-optimal demonstrations); DCTED and our PrObe get a slightly better result because they filter useful information from demonstrations.
Figure 14: PrObe’s failure prediction. In the Pick &\&& Place task, the rollout is terminated immediately if the mug is no longer on the table, which means the error detectors have only six frames to detect the error in this case. However, it took a while for our Probe’s pattern flow to cause enough change. By the last frame, ProObe was just about to increase its predicted probability, but it was too late.
Case study on PrObe’s failure prediction
We analyze the cases in which PrObe is poor at detecting behavior errors and visualize the results in Figure 14. As stated in the main paper, our PrObe requires a short period after the error occurs to allow the pattern flow to cause enough changes. However, in the Figure 14 case, the mug had fallen off the table, causing this rollout to be stopped immediately. Such a short time is insufficient for PrObe to reflect adequately, rendering it unable to (substantially) increase the prediction probability of behavioral errors at the last moment. Developing mechanisms for quick responsive pattern flow is one of our future research directions.
E.4 Pilot Study on Eeror Correction
To further examine the practicality of our AED task, we conducted a pilot study on error correction. In this experiment, the FSI policy is paused after the AED methods detect an error. Then, a correction policy from Wong et al. (2021), which resets the agent to a predefined safe pose, is applied. At last, the FSI policy continues to complete the task.
Table 7: Error correction results. Success Rate (SR) [$\uparrow$] is reported. The detection threshold is set as 0.9 for all compared methods. The values indicate the performance of the DCT policy without correction (first column) and its collaboration with the correction policy and four AED methods (remaining columns). PrObe is the only method that improves the performance, as it detects the error most accurately.

### v2

@@ Appendix A FSI Policy Details
We follow the main paper’s Section 3 to implement three state-of-the-art demonstration-conditioned (DC) policies, including NaiveDC[8], DCT[1], and SCAN[11], and monitor their behaviors.
@@
In the visual head, we added a fully-connected and a dropout layer after a single ResNet18 to mitigate overfitting of the data from base environments. Moreover, the input dimension of the actor module is determined by adding the dimensions of the outputs from the visual head and the task-embedding network. For the DCT and SCAN policies, the task-embedding’s dimension matches the visual head’s output. However, the task-embedding’s dimension in NaiveDC policy is twice the size.
Regarding the task-embedding network, NaiveDC concatenates the first and last frames of each demonstration Subsequently, it computes the average as the task-embedding, without involving any attention mechanism; DCT incorporates cross-demonstration attention to fuse demonstration features, followed by applying rollout-demonstration attention to compute the task-embedding; SCAN utilizes an attention mechanism to attend to each demonstration from the rollout, aiming to filter out uninformative frames. It then fuses attended frames to obtain the final task-embedding. Furthermore, we employ a standard LSTM in our implementation, differing from the setting described in [11].
@@
To optimize the policies, we utilize a RMSProp optimizer [49] with a learning rate of 1e-4 and a L2 regularizer with a weight of 1e-2. Each training epoch involves iterating through all base environments. Within each iteration, we sample five demonstrations and ten rollouts from the sampled base environment. The total number of epochs varies depending on the specific tasks and is specified in Table 4. All policy experiments are conducted on a Ubuntu 20.04 machine equipped with an Intel i9-9900K CPU, 64GB RAM, and a NVIDIA RTX 3090 24GB GPU.
@@ Appendix B PrObe Details
We propose the Pattern Observer (PrObe) to address the novel AED task by detecting behavior errors from the policy’s feature representations, which offers two advantages: (1) PrObe has a better understanding of the policy because the trained policy remains the same during both AED training and testing. This consistency enhances its effectiveness in identifying errors. (2) The additional task knowledge from the policy encoder aids in enhancing the error detection performance. We have described PrObe’s architecture and their usages in Section 5 of the main paper. Here, we introduce more PrObe’s design justification and training objectives.
PrObe’s design justification
The pattern extractor aims to extract observable patterns from embeddings in policy features. One potential approach is to leverage discrete encoding, as used in [50]. However, the required dimension of discrete embedding may vary and be sensitive to the evaluation tasks. To address variability, we leverage an alternative approach that operates in a continuous but sparse space. Additionally, the goal of the pattern flow generator is to capture the temporal information of current patterns. Due to the characteristics of the AED task, a standard LSTM model is better suited than a modern transformer. This is because policy rollouts are collected sequentially, and adjacent frames contain crucial information, especially when behavior errors occur. Finally, PrObe detects inconsistencies by comparing the pattern flow with the transformed task embedding. The contributions of PrObe’s components are verified and shown in Figure 12. The results demonstrate that our designed components significantly improve performance.
@@
With the objective $L_{pat}$, the pattern embeddings are expected to be sparse and easily observable of changes, benefiting the model in distinguishing behavior errors.
Next, we highlight that the time relationship should also be considered when applying contrastive learning to temporal-sequence data, in addition to determining whether the frames are normal or errors for the task. Thus, $L_{tem}$ is a novel temporal-aware triplet loss [40], and the temporal-aware margin $m_{t}$ in $L_{tem}$ will have different effects depending on the relationship between the anchor and the positive/negative sample, as depicted in Figure 7. The $L_{tem}$ can be calculated by:
@@
where $N$ is the number of sampled pairs, and the temporal-aware margin $m_{t}=m*(1.0-\alpha*(d_{ap}-d_{an}))$ will be enlarged or reduced considering the temporal distance of anchor, positive and negative samples. Among them, $m$ is the margin in the original triplet loss, and $d_{ap}$, $d_{an}$ are the clipped temporal distances between the anchor and positive sample and the anchor and negative sample, respectively. With $L_{tem}$, PrObe can efficiently perform contrastive learning without getting confused by blindly pulling temporally distant anchors and positive samples closer.
Lastly, the total loss $L_{total}$ is the combination of three objectives:
@@
method name policy training temporal DC-based output type ##\## of
dependent rollouts information parameters
SVDDED only successful distance 11.31M
TaskEmbED both probability 11.52M
LSTMED both probability 11.85M
DCTED both probability 14.49M
PrObe (ours) both probability 0.65M
@@ Appendix C Error Detector Details
  • SVDDED: A modified deep one-class SVM method [16] trained only with successful rollouts. It relies on measuring the distance between rollout embeddings to the center embedding to detect behavior errors, without considering the temporal information. We compute the center embedding by averaging demonstration features and minimize the embedding distance between rollout features and the center during training.
  • TaskEmbED: A single-frame baseline trained with successful and failed rollouts, which is a modification from the NaiveDC policy [8]. It concatenates and averages the first and last demonstration frames as the task-embedding. Then, it predicts the behavior errors conditioned on the current frame and task-embedding.
  • LSTMED: A deep LSTM model [45] predicts errors based solely on the current rollout. It is trained with successful and failed rollouts without access to demonstration data. It is expected to excel in identifying errors that occur at similar timings to those observed during training. However, it may struggle to recognize errors that deviate from the demonstrations.
  • DCTED: A baseline with temporal information derived from the DCT policy [1]. It is trained with successful and failed rollouts and incorporates with demonstration information. One notable distinction between DCTED and our PrObe lies in their policy dependencies. DCTED detects errors using its own encoder, which solely leverages visual information obtained within the novel environment. Conversely, PrObe learns within the policy feature space, incorporating additional task-related knowledge.
@@ Appendix D Evaluation Tasks
To obtain accurate frame-level error labels in the base environments and to create a simulation environment that closely resembles real-world scenarios, we determined that existing robot manipulation benchmarks/tasks [51, 41, 42, 52, 11, 43, 44] did not meet our requirements. Consequently, seven challenging FSI tasks containing 322 base and 153 novel environments are developed. Their general attributes are introduced in Section D.1 and Table 4; the detailed task descriptions and visualizations are provided in Section D.2 and Figure 8-9, respectively.
@@ D.1 General Task Attributes
  • Generalization: To present the characteristics of changeable visual signals in real scenes, we build dozens of novel environments for each task, comprising various targets and task-related or environmental distractors. Also, dozens of base environments are built as training resources, enhancing AED methods’ generalization ability. Letting them train in base environments with different domains can also be regarded as a domain randomization technique [9], preprocessing for future sim2real experiments.
  • Challenging: Three challenges, including multi-stage tasks, misaligned demonstrations, and different appearances between expert and agent, are included in our FSI tasks. Moreover, unlike existing benchmarks (e.g., RLBench [41]) attaching the objects to the robot while grasping, we turn on the gravity and friction during the whole process of simulation, so grasped objects may drop due to unsmooth movement.
  • Realistic: Our tasks support multi-source lighting, soft shadow, and complex object texture to make them closer to reality (cf. Figure 9). Existing benchmarks usually have no shadow (or only a hard shadow) and simple texture. Besides, all objects are are given reasonable weights. The aforementioned gravity and friction also make our missions more realistic.
@@ D.2 Task Descriptions
Description Fully close the drawer that was closed by the expert in demonstrations. Besides, the objects on the table will be randomly set in the designated area.
Success condition The target drawer is fully closed, and another drawer (distractor) must not be moved.
Behavior errors (1) Not fully close the target drawer; (2) Close the distractor, not the target drawer; (3) Close two drawers at the same time.
@@
Description Fully press the button which is randomly placed in a designated area.
Success condition The button is fully pressed.
Behavior errors (1) Not touching the button at all; (2) The button is not fully pressed.
@@
Description Pick and place the mug/cup into the bowl the expert placed in demonstrations.
Success condition The mug/cup is fully placed in the target bowl.
Behavior errors (1) Failed to pick up the mug/cup; (2) Failed to place it into the target bowl; (3) Misplaced the mug/cup into another bowl (distractor).
@@
Description Pick up and place a glass of water on the coaster the expert placed in demonstrations.
Success condition The glass of water is placed on the target coaster and no water is spilled.
Behavior errors (1) Failed to pick up the glass of water; (2) drips spilling; (3) Not placed on the target coaster.
@@
Description Pick up and place the object in front of the target bookshelf, and push it under the shelf.
Success condition The object is fully under the target shelf.
Behavior errors (1) Failed to pick up the object; (2) Losing object during moving; (3) Not fully placed under the target shelf.
@@
Description Pick up the magic cube/dice and place it into the storage box. Then, push the box until it is fully under the bookshelf.
Success condition The magic cube/dice is in the storage box, and the box is fully under the bookshelf.
Behavior errors (1) Failed to pick up the magic cube/dice. (2) Failed to place the magic cube/dice into the storage box. (3) The box is not fully under the bookshelf.
@@
Description Wait until the product box reaches the operating table, place the product in the box, and pick and place the box on the next conveyor belt.
Success condition The product is inside the box, and the box reaches the destination table.
Behavior errors (1) Failed to place the product into the box. (2) Failed to pick up the box. (3) Failed to place the box on the next conveyor belt.
@@ Appendix E Additional Experimental Results
E.1 Accuracy Function
Figure 10: Rules in the accuracy function $A(\cdot,\cdot)$ to determine the error prediction results.
As stated in the main paper, our control and understanding are diminished in novel environments. Consequently, during inference, we lack frame-level labels, only discerning success or failure at the end of the rollout. We adopt similar rules from [19] to define the accuracy function $A(\cdot,\cdot)$, as depicted in Figure 10. This function determines the outcome of error prediction at the sequence-level, which may not adequately reflect the accuracy of timing for error detection. Hence, we proceed to conduct the timing accuracy experiment (Figure 5 in the main paper) to address this concern.
@@ E.2 Main Experiment’s Details
Table 5: FSI policy performance. Success rate (SR) [$\uparrow$] and its standard deviation (STD) are reported. STD is calculated across novel environments, rather than multiple experiment rounds. This accounts for the variability in STDs, as the difficulty of novel environments within the same task can vary.
Close Drawer Press Button Pick &\&& Place Move Glass Cup
[…]
Organize Table Back to Box Factory Packing
[…]
@@
The policy performance is originally reported in Figure 4 of the main paper. Here, we provide a comprehensive overview in Table 5. Each policy conducts 20 executions (rollouts) for each novel environment to obtain the success rate (SR). Subsequently, the average success rate and its standard deviation (STD) are calculated across these SRs. Large STD values may arise due to the diversity in difficulty among novel environments. Factors such as variations in object size or challenging destination locations can lead to divergent performance outcomes for the policies. Additionally, these rollouts are collected for later AED inference.
In general, the NaiveDC policy achieves the best results in simple tasks because the task-embedding from the concatenation of the first and last two frames provides sufficient information. However, its performance rapidly declines as task difficulty increases. Besides, the DCT policy excels in tasks where the misalignment between demonstrations is less severe, as it fuses the task-embeddings in the temporal dimension. Lastly, the SCAN policy outperforms in tasks with high temporal variability because its attention mechanism filters out uninformative demonstration frames effectively.
Main experiment’s ROC curves
We have reported detailed numbers from the main experiment in Figure 4 of the main paper. Here, Figure 15 (on page 25) presents the ROC curves of all tasks. The numbers reported in Figure 4 represent the area under the curves. To reiterate, our proposed PrObe achieves the best performance on both metrics and can handle the various behaviors exhibited by different policies.
@@ E.3 Ablation Study
Rollout augmentation
This experiment aims to validate whether rollout augmentation (RA) enhances the collected agent rollout diversity and, consequently, improves the robustness of error detectors. Figure 11 presents the results of all error detection methods (trained with and without RA) monitoring policies solving the Pick &\&& Place task, yielding the following findings: (1) RA has a minor or even negative impact on SVDDED, as expected. Since SVDDED is exclusively trained on single frames from successful rollouts, the removal of frames by RA decreases overall frame diversity. (2) RA proves beneficial for methods incorporating temporal information (LSTMED, DCTED, and PrObe), particularly when the FSI policy performs better (SCAN case). This is because error detectors cannot casually trigger an error, or they risk making a false-positive prediction. The results indicate that methods trained with RA generate fewer false-positive predictions, showcasing improved robustness to various policy behaviors.
@@
PrObe’s design verification
Figure 12 illustrates the performance contribution of each component of PrObe, observed while monitoring the SCAN policy solving the Pick & Plcae task. The naive model (first column) removes the pattern extractor and utilizes a FC layer followed by an IN layer to transform history embeddings. Clearly, the designed components and objectives enhance performance, particularly with the addition of the pattern extractor into the model. We attribute this improvement to the fact that the extracted embedding patterns by our PrObe are more informative and discernible.
@@
Performance stability
Our work involves training and testing all policies and AED methods to produce results for an FSI task. Additionally, each task comprises numerous base and novel environments. These factors make it time-consuming and impractical to include error bars in the main experiment (Figure 4 of the main paper). Based on our experience, generating the results of the main experiment once would require over 150 hours.
Therefore, we selected the most representative Pick &\&& Place task among our FSI tasks and conducted multiple training and inference iterations for all AED methods. We present the AUROC, AUPRC, and their standard deviations (STD) averaged over multiple experimental rounds (with five random seeds) in Figure 13. From the results, our proposed PrObe not only achieves the best performance in detecting behavior errors across all policies but also exhibits the most stable performance among all AED methods, with the smallest STD values.
Figure 14: PrObe’s failure prediction. In the Pick &\&& Place task, the rollout is terminated immediately if the mug is no longer on the table, giving the error detectors only six frames to detect the error in this case. However, it took a while for our Probe’s pattern flow to induce enough change. By the last frame, PrObe was just beginning to increase its predicted probability, but it was too late.
PrObe’s failure prediction
We analyze cases in which PrObe is ineffective at detecting behavior errors and visualize the results in Figure 14. As stated in the main paper, our PrObe requires a short period after the error occurs to allow the pattern flow to induce enough changes. However, in the case depicted in Figure14, the mug had fallen off the table, resulting in the immediate termination of the rollout. Such a short duration is insufficient for PrObe to adequately reflect, thereby rendering it unable to substantially increase the prediction probability of behavioral errors at the last moment. Developing mechanisms for quick responsive pattern flow is one of our future research directions.
Table 6: Demonstration quality experiment. AUPRC [$\uparrow$] is reported. This experiment verifies the influence of demonstration quality on both the FSI policy and AED methods. PrObe is the only method that can achieve higher performance when receiving sub-optimal demonstrations, as the FSI policies raise more failures.
@@
[…]
@@
We verify the influence of demonstration quality on both the FSI policy and AED methods, specifically testing it on the Press Button task. We collected sub-optimal demonstrations that failed to press the button on the first trial but successfully pressed it on the second try. Then, the policies and AED methods used these demonstrations to perform the task.
We expected the policy performance to decrease as the demonstration quality decreases. Therefore, AED methods should achieve a higher AUPRC score if they are not affected by the quality changes of demonstrations, as it becomes easier to detect errors. However, from Table 6, we observe that not all AED methods obtain higher scores, indicating that they are also affected by the decrease in quality.
Notably, our proposed PrObe is the only method that achieves better results for all cases when comparing the first and second rows, and the first and third rows for each policy. Additionally, an interesting observation can be found in the case of NaiveDC policy (comparing all optimal vs. all sub-optimal): LSTMED and TaskEmbED obtained similar AUPRC scores, as they do not access the demonstration information directly; SVDDED suffered in this situation since it averages every demonstration frames (including those sub-optimal movements); DCTED and our PrObe achieved slightly better results because they filter useful information from the demonstrations.
Table 7: Viewpoint changes. AUROC [$\uparrow$] is reported. We study the impact of viewpoint changes on AED methods’ performance. The camera is slightly shifted away from the table during inference, which makes detecting errors more difficult. Nevertheless, our PrObe is the least affected method.
Viewpoint SVDDED TaskEmbED LSTMED DCTED PrObe
original 0.5694 0.5465 0.4929 0.7501 0.7498
shift 0.4821 0.4959 0.4667 0.7056 0.7225
[…]
Viewpoint changes
[…]
From Table 7, although the policy yields more behavior errors after camera shifted, the performance of all AED methods decreases due to the increased difficulty of detecting errors. Nevertheless, our PrObe is the least affected method, which demonstrates its superiority against other approaches.
E.4 Pilot Study on Error Correction
Table 8: Error correction results. Success Rate (SR) [$\uparrow$] is reported. The detection threshold is set as 0.9 for all methods. The values indicate the performance of the DCT policy without correction (first column) and its collaboration with the correction policy and four AED methods (remaining columns). PrObe is the only method that improves the performance, as it detects the error most accurately.